# Uncertainty Calibration for Counterfactual Propensity Estimation in Recommendation

### v2

@@ V-B Recommendation with Selection Bias
Doubly Robust (DR) estimators [69, 10] simultaneously account for imputation errors and propensities to reduce variance in IPS. Recent improvements include asymmetric tri-training [70], information theory considerations [71], adversarial training [72], enhanced doubly robust estimators [1], knowledge distillation [17], bias-variance trade-off [2], and multi-task learning [3].
@@ V-C Uncertainty Calibration and Quantification
Beyond Platt scaling, temperature scaling has been proposed for uncertainty calibration in multi-class classification [12]. Platt scaling and temperature scaling both assume a Gaussian distribution. For distributions that are richer and more skewed, Beta calibration is another effective method [32].
In addition to parametric methods that assume distributional assumptions, non-parametric techniques can also be considered. These include histogram binning [76] and isotonic regression [77].
@@
From a Bayesian generative model perspective, this paper focuses on approximate methods like MC Dropout and Deep Ensembles for uncertainty quantification in IPS. While Gaussian processes, Bayesian neural networks, and other probabilistic graphical models can also quantify uncertainty in IPS [78], practical approximate inference on large-scale datasets poses significant challenges [79, 80].
@@
    Deep ensembles, while providing robust uncertainty estimates, are computationally expensive due to the necessity of training multiple models [35]. Hence the BatchEnsembles method can be used to reduce the overall computation cost as detailed in Sec. III-B3 [44].
@@
VI Conclusions
In this paper, we first propose the use of Expected Calibration Error (ECE) to assess the quality of propensity scores and highlight their uncertainty miscalibration in recommendation scenarios with data missing not at random. We introduce uncertainty calibration methods for propensity score estimation and compare three calibration techniques for this purpose. Our theoretical analysis demonstrates that calibrated Inverse Propensity Scores (IPS) exhibit reduced bias.
Experimental results on two representative datasets, Coat Shopping and Yahoo! R3, confirm that calibrated inverse propensities significantly enhance recommendation accuracy. These findings underscore the importance of addressing propensity score uncertainty in mitigating bias and improving recommendation systems.

### v3

@@ V-B Recommendation with Selection Bias
Doubly Robust (DR) estimators [74, 10] simultaneously account for imputation errors and propensities to reduce variance in IPS. Recent improvements include asymmetric tri-training [75], information theory considerations [76], adversarial training [77], enhanced doubly robust estimators [1], knowledge distillation [17], bias-variance trade-off [2], and multi-task learning [3]. DR-V2 [53] proposes balanced-mean-squared-error metric for joint propensity and CVR estimation. [78]presents a novel combinational joint learning framework that simultaneously learns unbiased user-item relevance and propensity estimation to improve the accuracy of implicit recommender systems.[79] leverages both user and item perspectives to estimate propensity scores, addressing biases in sequential recommendation systems. [80] introduces DDPO, a framework that mitigates sample selection bias in post-click conversion rate estimation by optimizing models with both clicked and unclicked samples in the impression space.
@@ V-C Uncertainty Calibration and Quantification
Beyond Platt scaling, temperature scaling has been proposed for uncertainty calibration in multi-class classification [12]. Platt scaling and temperature scaling both assume a Gaussian distribution. For distributions that are richer and more skewed, Beta calibration is another effective method [32]. In addition to parametric methods that assume distributional assumptions, non-parametric techniques can also be considered. These include histogram binning [84] and isotonic regression [85].
@@
From a Bayesian generative model perspective, this paper focuses on approximate methods like MC Dropout and Deep Ensembles for uncertainty quantification in IPS. While Gaussian processes, Bayesian neural networks, and other probabilistic graphical models can also quantify uncertainty in IPS [86, 87], practical approximate inference on large-scale datasets poses significant challenges [88, 89]. Different methods can excel in specific scenarios, but each has its limitations. A summary of common drawbacks for various approaches is deferred to Section V in the Supplemental Materials.
VI Conclusions
This paper introduces Expected Calibration Error (ECE) as a novel metric to evaluate the reliability of propensity scores, shedding light on the prevalent issue of uncertainty miscalibration in recommendation systems where data is missing not at random (MNAR). To address this challenge, we propose uncertainty calibration techniques for propensity score estimation and systematically compare three calibration approaches.
Through theoretical analysis, we demonstrate that calibrated Inverse Propensity Scores (IPS) reduce bias, leading to more reliable debiasing in recommendation tasks. Extensive experiments on three benchmark datasets, Coat Shopping, Yahoo! R3 and KuaiRand, validate the effectiveness of our methods, showing that calibrated propensity scores significantly enhance recommendation accuracy.
These findings emphasize the critical role of addressing propensity score miscalibration in improving both bias mitigation and the overall performance of recommendation systems. This work highlights a promising direction for incorporating uncertainty calibration to ensure more robust and fair recommendations.
Supplemental Materials for
Uncertainty Calibration for Counterfactual Propensity Estimation in Recommendation
VIII Summary of Main Symbols
Table VII describes the main symbols used in this paper.
Symbol Description
$\mathcal{U}$ Users set
$\mathcal{I}$ Items set
$\mathbf{R}$ conversion matrix
$\mathbf{\hat{R}}$ predicted conversion matrix
$\mathcal{O}$ click label matrix
$\mathbf{R^{o}}$ observed conversion matrix
$\mathcal{D}$ user-item pairs space
$\mathbf{E}$ prediction error matrix
$\mathbf{\hat{E}}$ imputed error matrix
$\mathcal{P}$ propensity scores matrix
$\mathcal{\hat{P}}$ estimated propensity scores matrix
$g_{\phi}$ propensity estimation model
$f_{\theta}$ CVR prediction model
TABLE VII: The summary of main symbols used in this paper.
Figure 3: The relationship between ECE and recommendation metrics.
IX Model implementation
We implement all models with Pytorch and optimize them with adam optimizer. We first determine the hyper-parameters for NeuMF(backbone) based on grid search, and the search range for the embedding size, batch size, learning rate, L2 regularization coefficient are set as 16, 32, 64, 128, 256, 256, 512, 1024, 2048, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2 and 1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3 respectively. The best configuration for each mothod is determined based on the ranking performance on the validation set. The search results is as follows: the embedding size, batch size, learning rate, dropout rate and L2 regularization coefficient are set to 64, 1024, 0.001, 0.2 and 1e-4 respectively. The structure of the MLP layers in NeuMF is set to [64, 32, 16]. Other models, including IPS, DR-JL, MRDR, CDR, GPL, ESCM2 and DR-V2, are all built upon NeuMF. They use the same hyper-parameter settings as the baseline NeuMF for common hyper-parameters.
For evaluating recommendation results, we employ three metrics: AUC, discount cumulative gain (DCG) and Recall [1]. DCG and Recall are defined as:
${\rm DCG}(K)=\sum_{k=1}^{K}\frac{Rel_{k}}{\log_{2}(k+1)},$ (24)
${\rm Recall}(K)=\sum_{k=1}^{K}Rel_{k},$ (25)
X The relationship between calibration error and recommendation results.
In this section, we have conducted a detailed analysis of how improved calibration impacts recommendation metrics. Specifically, we utilized Platt scaling as a post-hoc calibration method to adjust propensity scores. Platt scaling optimizes the negative log-likelihood (NLL) loss using the LGFBS optimizer, where a reduction in NLL is accompanied by a corresponding decrease in Expected Calibration Error (ECE).
To demonstrate the relationship between ECE and recommendation performance, we set checkpoints every 10 epochs during training, saving the propensity scores, ECE, and NLL values at each point. We then trained an Inverse Propensity Scoring (IPS) model on the Coat dataset using these saved propensity scores to evaluate recommendation metrics, including AUC, DCG, and Recall. The results, as illustrated in the figure 3, clearly show a strong negative correlation between ECE and recommendation performance: as ECE decreases, metrics such as AUC, DCG, and Recall exhibit significant improvement.
This empirical evidence supports our claim that better-calibrated propensity scores (lower ECE) lead to superior recommendation outcomes, thereby providing a clearer and more compelling link between calibration quality and CVR prediction enhancement. Hence, our findings affirm that reducing ECE improves IPS predictions and overall recommendation performance.
XI Calibration Curve and Propensity Histogram of Calibrated Propensity scores on the Yahoo! R3 Shopping Dataset
(a) Calibration Curve
(b) propensity scores Histogram
XII Shortcomings of Each Calibration Method
@@
    Deep ensembles, while providing robust uncertainty estimates, are computationally expensive due to the necessity of training multiple models [35]. Hence the BatchEnsembles method can be used to reduce the overall computation cost as detailed in [48].
@@
  4. 4.
    Dual Focal Loss effectively addresses class imbalance and hard-to-classify examples but has notable shortcomings. Dual Focal Loss can overfit noisy data by overly focusing on mislabeled or ambiguous examples. It requires tuning of hyperparameters like the focusing factor, adding complexity to training [47].